Show on the dimensional Analysis that v^2 - u^2 = 2as

1
Expert's answer
2022-01-10T09:10:31-0500

Dimension of a speed

"[v]=[u]=\rm m/s=LT^{-1}"

Dimension of an acceleration

"[a]=\rm m/s^2=LT^{-2}"

Dimension of a distance

"[s]=\rm m=L"

Let's check an equation

"v^2 - u^2 = 2as"

We get

"[v^2]-[u^2]=[a][s]"

"\rm (LT^{-1})^2-(LT^{-1})^2=(LT^{-2})L"

"\rm L^2T^{-2}=L^2T^{-2}"

We get an identity, so equation is dimensionally correct.